The cost of 6 meters of ribbon is $15. How much ribbon could you buy for $20


1
Expert's answer
2022-01-12T09:00:30-0500



Cost(C) "\propto" Length (L)


"\implies\>C=KL" where "K" is a constant

When "C=15"

"L=6"


"\implies\>15=K(6)"

"K=\frac{15}{6}"


"C=\frac{15}{6}L"


"\implies\>L=\frac{6}{15}C"


When "C=20"

"L=\frac{6}{15}×20"


"=8\>meters"